216 x 3 + 1/ 125

216 x3  + 1/ 125 

Grade:9

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given 216x3+1/125
It can be simplified as (6x)3+(1/5)3............(1)
we know a3+b3=(a+b)(a2-ab+b2)
so by expanding equation (1) we get
where a=6x and b=1/5
so (6x)3+(1/5)3=(6x+1/5)(36x2-(6x/5)+1/25)